The solar hot water system market in Australia receives a 25% price subsidy (corrective subsidy) from the government on each unit sold. The product sells at a market price (i.e. the subsidised price) of $1,875. A firm proposes a project to supply 50,000 additional systems per year. The annual cost of the project (operating cost plus annual equivalent capital cost) at both market and efficiency prices is $95 million. Calculate the net annual benefit of the project:

(a) AT MARKET PRICE

(b) AT EFFICIENCY PRICE
Answer #1

If the product is sold at market price of $1,875 then, revenue= $1875* 50,000= 93,750,000

Benefit= Revenue- Cost= 93,750,000- 95,000,000= (1,250,000) loss

If the product is sold at market price + price of subsidy (efficiency price)=

Revenue= 1875*700/75= 2500

Revenue= $2500*50000= $125 million

Benefit= $125 million - $95 million= $35 million.